Apply Charpit's method to solve the following differential equation:
p(1+q^2)=q(z-a)
1
Expert's answer
2020-12-01T06:26:25-0500
"Solution"

Here, given equation is:


"p(1+q^2)=q(z-a)\\
\therefore p+pq^2)=qz-qa\\
 \implies p+pq^2-qz+qa=0---(1)"

Therefore the Charpit's auxiliary equation is


"\frac{dp}{\frac{df}{dx}+p\frac{df}{dz}}=\frac{dq}{\frac{df}{dy}+q\frac{df}{dz}}=\frac{dz}{-p\frac{df}{dp}-q\frac{df}{dq}}=\frac{dx}{-\frac{df}{dp}}=\frac{dy}{-\frac{df}{dq}}=\frac{dF}{0}"

or


"\frac{dp}{0+p(-q)}=\frac{dq}{0+q(-q)}=\frac{dz}{-p-pq^2}=\frac{dx}{-1-q^2}=\frac{dy}{-2pq-z}=\frac{dF}{0}"

or


"\frac{dp}{-pq}=\frac{dq}{-q^2}=\frac{dz}{-p-pq^2}=\frac{dx}{-1-q^2}=\frac{dy}{-2pq-z}=\frac{dF}{0}"

So that


"\implies \frac{dp-dq}{-pq+q^2}=\frac{dy-dx}{-2pq-z+1+q^2}\\
\implies \frac{dp-dq}{-pq+q^2}=\frac{dy-dx}{-pq+q^2}\\
\implies dp-dq = dy-dx\\
\therefore dp+dx=dy+dq\\"

Integrating, we get


"\implies p+x=y+q\\
\implies p=y-x+q\\"

From equation (1), we get


"p+pq^2-qz+qa=0 \implies p=pq^2-qz+qa\\
 q(pq-z-a)=p\\
q=p\ or\\
q=\frac{z-a}{p}"

Using the value "p" and "q" in "dz=pdx+qdy"


"\implies dz=(y-x+q)dx+(\frac{z-a}{p})dy\\
\implies \frac{dz}{z}=(y-x+q)dx+(\frac{z-a}{p})dy"


Integrating, we get


"log\ z=yx+qx-\frac{x^2}{2}+\frac{y-ya}{py}"